find the equation of the line containing points ( 0, -3) and ( -12, -5)
Answer by rfer(16322)

m=(-5--3)/(-12-0)
m=-2/-12
m=1/6
----------------
y=mx+b
-5=1/6*-12+b
-5=-2+b
-3=b
----------
y=1/6x-3